how much antifeeze should be added to 300 gallons of 30% antifreeze solution to raise the antifreeze solution to 40%?
Answer by ptaylor(1894)

Let x=amount of pure antifreeze that needs to be added
Now we know that the amount of pure antifreeze in the quantity that was added (x) plus the amount of pure antifreeze in the 300 gal (0.30*300) has to equal the amount of pure antifreeze in the final mixture(0.40(300+x)). So our equation to solve is:
x+0.30*300=0.40(300+x) get rid of parens (distributive law)
x+90=120+0.40x subtract 0.40x and also 90 from each side
x+90-90-0.40x=120-90+0.40x-0.40x collect like terms
0.60x=30 divide each side by 0.60
x=50 gal-------------------------------amount of pure antifreeze needed
CK
50+90=0.40(350)
140=140

Hope this helps------ptaylor